OpenStudy (anonymous):

Compute: (4 + 9) − 3(6) ÷ 2 + 8

OpenStudy (anonymous):

12

OpenStudy (anonymous):

To solve this, remember order of operations: Parenthesis, Exponents, Multiplication/Division, then Addition/Subtraction, which you can remember by the mnemonic "Please Excuse My Dear Aunt Sally" (apparently Aunt Sally couldn't remember the order of operations. Don't be like Aunt Sally.). So: Perform the operations inside parenthesis (4+9) first. That gives us: \[13-3(6)\div2+8\] Then multiplication of 3(6) gives us: \[13-18\div2+8\] Then division (use the two immediate terms on either side of the division sign 18 / 2) gives us: \[13-9+8\] Finally, addition and subtraction gives us: \[12\] :)

OpenStudy (anonymous):

Start with do parentese's first, then multiplication and division. do 4+9=13, 3x6=18, then 18/2=9. So, you're equation will look like this, so far: 13-9+8= .... So just do, 13-9=4+8=12 So, you're final answer is 12!